# Living with Data: Exploring Physicalization Approaches to Sedentary Behavior Intervention for Older Adults in Everyday Life

### v2

@@ 4.1. RQ1: Mitigating Data Anxiety through Atmospheric Abstraction
Through observations and interviews, we found that many participants struggled with smartphones and avoided “one more digital device”. We also found cushions repeatedly appeared as mundane but central objects in long sitting episodes (living rooms, bedrooms), which have opportunities to highlight the intelligence of a particular data-physicalized thing that can sustain intimate entanglements in everyday settings align with previous work (jansen2015opportunities; frauenberger2019entanglement; Yu2025). Embedding sensing into the cushion promised low additional burden and minimal change to their physical environment.
4.1.2. Creating an Meaning-Making through Atmosphere (Phase 2).
Because connecting the choice of material and embodiment to interactions can help tailor a physicalization to a particular target audience (li2025crafting; miao2025physlens), we refined the concept into a two-part hybrid system: a sensing cushion plus an external display.
Initially we tried numeric and bar chart visualizations, assuming they would support clear self-monitoring. But in the poster feedback from both older adults and experts revealed strong negative reactions: graphs felt clinical, judgmental, and reminiscent of hospitals, raising concerns about data anxiety and long‑term engagement. This pushed us towards ambient, non‑numeric feedback using a small, furniture‑like object inspired by (Bae2022; suzuki2017fluxmarker).
@@ 4.1.3. Leveraging Abstraction and Feelings of Companionship for Meaning Making (Phase 3)
This sense of companionship was achieved by changing the room’s feeling. Participants perceived the device not as an information display but as felt it could integrate part of their home’s ambiance, providing a canvas upon which they could construct personal significance. P9_\_Phase3 said “It is not data, it is like an atmosphere, like when the sun sets and the room gets warm and golden, you know the day is ending. This light is like that. It doesn’t count at me. it just changes the feeling of the room. This change in lighting is very romantic … a very gentle way. Its a very poetic way of knowing time has passed.”. By comparing the light to a sunset, the participant was not merely having an experience, he was actively interpreting the abstract data by connecting it to a familiar, natural phenomenon. This act of drawing an analogy transformed the device from an information display into a part of the room’s narrative. This demonstrates that for this user group, data accessibility is contingent not only on simplification but also on its integration into a form that allows for personal meaning-making.
@@ 5. Discussion
Our study aimed to explore how to make sedentary data accessible for older adults and how this method differs from traditional reminders. The results show that the key to accessibility is not to simplify quantitative data but to transform it into an atmospheric abstraction (10.1145/3706598.3714302; Heijboer2016Peripheral). Participants reported experiencing the light as gentle, non-clinical, friendly, and as an ordinary lamp that did not make them feel judged or subjected to a medical check-up. This is in line with prior work that argues for supportive instead of paternalistic health technologies for older adults (Yu2025; voinea2024paternalistic). In this sense, our findings showing that supportive and atmospheric abstractions can be maintained in everyday in-situ use, rather than only in short-term or laboratory settings. Herein, we reflect on our findings and describe their implications for future interventions that support older adults living, and interacting with, their personal data.
@@ 5.1. Rethinking Forms of Physical Data Representations for Older Adults
In light of our results, it is necessary to shift the model of interaction from a one of command and compliance to one of suggestion and negotiation, which better protects a user’s sense of autonomy. Because the system did not provide precise numbers, it avoided generating feelings of being judged and monitored. Participants thus tended to perceive the device as a gentle and companion-like presence (Liu2023VoiceAssistantsCompanionship). Such positive emotional connections are difficult for traditional health tracking devices to establish even though they can make health interventions feel more humane and could increase the likelihood they are accepted by users in the long term.
The ambiguity and gentleness of such designs may have potential negative impacts as well. If a design is too focused on being non-intrusive, it may struggle to achieve its original health intervention goals (Frissen2024ProgressFeedback; Ntsweng2025AALLessons; Lee2023AmbientToDisruptive). When users ignore a reminder, it is also difficult to determine if this was due to an empowered, autonomous choice or if a design had failed. Rather than trying to find a perfect balance between clarity and ambiguity it might be more prudent to actively manage this uncertainty and explicitly give final control to the user (10.1145/3532106.3533471). Our mute button allowed users to negotiate with the system and acknowledge a reminder while choosing to act on it later. Allowing users to negotiate, delay, or selectively decline prompts may help reduce feelings of being controlled or judged and may lower reactance and data-related anxiety (thudt2018self; ramesh2024exploratory). It can also help preserve a sense of autonomy and dignity, which prior work has identified as a central concern in the design of technologies for ageing populations (li2025crafting; lupton2017feeling). In line with the exploratory, in-situ nature of our RtD process, we suggest that such negotiated engagement should be seen as an informative outcome that complements more conventional measures of short-term effectiveness.
@@ 5.2. Leveraging Research Through Design for Exploratory Digital-to-Physical Translations
This experience thus suggests that for this class of problem, RtD can act as a form of experiential inquiry. As our central question was phenomenal rather than technical (i.e., what does it feel like to live with an ambient representation of your own data?), we needed to built artifacts and have end users live with them (10.1145/291224.291235; 10.5555/1841406). Through this process, we uncovered the direct conflict between a desire for calm ambiance and the need for a clear, actionable alert (Lee2023AmbientToDisruptive; Pousman2006AmbientTaxonomy). Our artifacts became physical arguments, each one testing a different stance on this conflict.
@@ 5.3. Design Implications
These implications outline a design direction for future tangible health technologies that shifts the emphasis from monitoring and compliance to empowerment, allowing older adults to maintain dignity and autonomy while data are woven into their everyday life. In the context of health support, enabling older adults to sometimes delay, negotiate, or selectively ignore prompts, will help reduce reactance, avoid data anxiety, and make it more likely that the intervention will be lived with over time.
@@ 5.4. Limitations and Future Work
While providing novel insights, our research does have several limitations. First, the one day deployment of the CushLumi system during Phase 3 prevented us from understanding the potential long-term effects of the intervention, such as what occurs when its novelty wore off or what would occur if it was utilized outside a controlled, in-home setting. Our future plans include a longitudinal, in-situ study lasting several months to examine whether and how the atmospheric abstraction supports sustained changes in sedentary patterns and ongoing engagement over time. For instance, we are interested in understanding the degree to which atmospheric abstraction can lead to sustained behaviour change, and how users’ interpretations of the ambient light shift with prolonged use.
The small sample size of five older adults in Phase 3, also limits the generalizability of our findings to the broader older adult population and to other cultural or geographic contexts. As our research was conducted with older adults in an Australian, Western context future work must examine how ambient data physicalizations are perceived in other cultural settings, including non-Western contexts that may have different norms around aging, health, family dynamics, and family care.
Objective studies of sedentary behavior also typically rely on accelerometers within wearable devices to measure sedentary time (healy2008objectively; healy2008breaks). Our cushion-based sensing plays a similar role, but within a piece of decor. We did consider the use of load cells, pressure mats, and camera based sensing to measure sedentary time, however, load cells and mats required more rigid structures and careful calibration, while cameras raised privacy concerns. Flexible sensors offered a good trade-off between simplicity, comfort, and ease of replication. It would, however, be worthwhile to explore other sensing methods and form factors in the future.
Architecturally, the CushLumi system was also designed so that participants’ own sensor data could serve as a basis for secondary modelling and personalisation. This adaptive layer was not activated in the present study but will be explored in future work to better understand the wellbeing data. Furthermore, we plan to iterate on the CushLumi system’s design by exploring different materials, forms, and sensory modalities (e.g., subtle temperature changes) to expand the vocabulary of atmospheric physicalizations that are possible.
@@ 6. Conclusion
Using a three phase Research through Design methodology, our research explored how sedentary data could be translated into an atmospheric, ambient physicalization for older adults. By offering gentle, suggestive guidance rather than intrusive commands, participants’ data anxiety was mitigated and older adults felt empowered to be less sedentary. Our CushLumi System prototype fostered positive, relational engagement with one’s personal health data, marking a shift from merely informing users to enabling them to live reflectively with their data.
Appendix
@@ Appendix A Design Explorations and Rationale via RtD
We initially explored the use of a smartphone application that aimed to deliver sedentary behavior feedback using conventional interaction methods and visualizations. During Phase 1, however, participants had an aversion to complex interfaces, small text, and the cognitive load of managing yet another app. For example, one participant remarked, “If I can choose, I don’t want to use smart phone, it has too many functions” (P8). Participants also explicitly stated that they wanted to avoid reminders that felt like nagging. This feedback, in addition to prior work that identified that longer sedentary periods are associated with higher clustered metabolic risk and all‑cause and cardiovascular mortality (owen2010sedentary; healy2008objectively; healy2008breaks) and recommendations that emphasized interrupting sitting every 30 to 60 minutes (owen2010sedentary; healy2008breaks; martin2015interventions; prince2014comparison), led to the exploration of interventions that could be embedded within existing physical environments and daily routines to discourage sedentary behavior.
@@
Building on the feedback obtained during Phase 2, we then designed and implemented the CushLumi System (Figure 2). Our design included a cushion and a light because prior ambient display work leveraged light for slowly changing, low priority information(wisneski1998ambient; davis2017effects; mclaughlin2020designing; feine2025integrating) and research on aging in place environments highlighted furniture (e.g.,lighting) as a key interior design factor for supporting sleep, mood and daily functioning in older adults (Yu2025).
To make the connection between the formative findings and the final form factors explicit, we summarize our design explorations across the three RtD phases in LABEL:tab:design_rationale. For each phase, we articulate the design goals, the rationale grounded in user and expert feedback, the configuration we chose to pursue next, and representative participant quotes that directly motivated these choices.
@@
cushion as form factor
Understand older adults’ sedentary routines and attitudes to technology; explore initial ways to surface sitting time without disrupting everyday practices. Through observations and interviews, we found that many participants struggled with smartphones and avoided “one more digital device”. We also found Cushions repeatedly appeared as mundane but central objects in long sitting episodes (living rooms, bedrooms), which have opportunities to highlight the intelligence of a particular data-physicalized thing that can sustain intimate entanglements in everyday settings align with previous work (jansen2015opportunities; frauenberger2019entanglement; Yu2025). Embedding sensing into the cushion promised low additional burden and minimal change to their physical environment. “If I can choose, I don’t want to use smart phone, it has too many functions.” (P8) [On cushions] “This is something I always use when I sit here.” (P1, field notes)
@@
Clarify how sedentary data should be represented; test reactions to explicit quantitative displays; maintain comfort by separating sensing from display. Because connecting the choice of material and embodiment to interactions can help tailor a physicalization to a particular target audience (li2025crafting; miao2025physlens), we refined the concept into a two-part hybrid system: a sensing cushion plus an external display. Initially we tried numeric and bar-chart visualizations, assuming they would support clear self-monitoring. But in the poster feedback from both older adults and experts revealed strong negative reactions: graphs felt clinical, judgmental, and reminiscent of hospitals, raising concerns about data anxiety and long‑term engagement. This pushed us towards ambient, non‑numeric feedback using a small, furniture‑like object inspired by (Bae2022; suzuki2017fluxmarker). “It reminded me of how I felt when I needed to go to the hospital.” (P15, on bar charts) “Corny”, “boring”, and “annoying”. (P15, on group discussion) “I feel a bit oppressed by it.” (P13, on explicit numeric feedback) It’s not just telling me good job for standing, it’s reminding me of something I love.(P11, on explicit numeric feedback)
@@
Realise a fully functional system that embodies an “ambient, gentle” data narrative; keep sensing invisible while offering legible but non‑intrusive feedback; support user agency and positive associations with movement. In order to build the composite relation with a data-physicalized thing in terms of non-neutrality, mutual shaping, and shared intentionalities, we implemented the CushLumi system: the cushion stays visually unchanged and embeds flex sensors plus a microcontroller to detect sitting (zhong2025investigating). We aim to build and deployed the technological data-physicalized thing as it has intentionality in assembling the cohabiting human beings and other things in the situated context of everyday (wakkary2021things). At the same time, we also hope that through the cube’s form design, we can offer target users more activity suggestions after standing, and motivate users to stand up proactively through intuitive visual cues (anderson2011seductive). A small wooden “ambient light” box sits on nearby furniture and gradually brightens as sitting accumulates; when the person stands, it shifts to a pleasant colour shining through laser‑cut patterns of everyday activities (e.g., stretching, gardening), gently suggesting alternatives to prolonged sitting. A touch‑sensitive area lets users mute feedback, reinforcing a sense of control. Participants often described the box as “a little lamp” or “a decoration”, not as a life countdown reminder. “I didn’t feel pushed. I just thought, ‘Maybe I’ll water the plants.’ And I did.” (P4, deployment interview)
Appendix B CushLumi System Implementation Details
The Sensing Cushion contained two Spectra Symbol 4.5 inch flexible sensors that were embedded in the foam of a 30 cm × 60 cm foam cushion. The sensors formed a shallow V pattern that was intended to roughly align with a user’s thighs and pelvis. Each flex sensor was sampled by a Particle Photon microcontroller at 10 Hz. The microcontroller read both analog channels, detected bending due to body weight, and aggregated readings over a 3 second window with hysteresis.
The Ambient Light was a small wooden box that featured laser-cut patterns of positive activities (e.g., stretching, gardening) on its faces, giving it the appearance of a tabletop lamp. The lamp measured approximately 8 cm × 8 cm × 8 cm and was powered by a single-cell LiPo battery, allowing it to be freely placed on side tables or shelves. The box contained 16 addressable WS2812 RGB LED Circle that glowed depending on one’s sitting time (i.e., 0 - 20 minutes -¿ no intervention needed so the LEDs were warm white, at 20 - 30 percent brightness; 20 - 40 minutes -¿ gentle prompting where the hue linearly interpolated from warm white to orange/red as a function of the elapsed minutes, with the brightness increasing moderately over time; 40+ minutes -¿ standing is recommended so the hue shifted toward saturated red, with the brightness increased to 70 to 80 percent, with smaller additional increases if the user did not stand up). In line with work showing that even replacing small amounts of sedentary time with light intensity movement can benefit metabolic health (healy2008objectively; healy2008breaks), the mapping was intended to gently nudge users toward more frequent, small breaks instead of signaling a sudden point where it was already too late to see benefits of standing. The intensity levels conservatively operationalized the 30 to 60 minute recommendations reported in sedentary behavior research (owen2010sedentary; healy2008breaks; dempsey2020new). When a user stood up, a short celebratory animation was played (i.e., the lamp cycled through rainbow hues for 2 to 3 seconds). The lamp’s LEDs were updated at a frequency of 1 Hz. A touch-sensitive area on the top served as a Mute button, providing direct control over the feedback and reinforcing the user’s sense of agency. Five CushLumi Systems were built for use in Phase 3.

### v3

@@ 4. Findings
Our analysis reveals how shifting from explicit numerical tracking to ambient physicalization fundamentally alters older adults’ relationship with their health data. We structure these findings to illustrate how atmospheric abstraction can mitigate data anxiety (RQ1) while fostering a sense of user agency through suggestive guidance (RQ2).
@@ 4.1. RQ1: Mitigating Data Anxiety through Atmospheric Abstraction
Through observations and interviews, we found that many participants struggled with smartphones and avoided “one more digital device”. We also found that cushions repeatedly appeared as mundane but central objects in long sitting episodes (living rooms, bedrooms), which have opportunities to highlight the intelligence of a particular data-physicalized thing that can sustain intimate entanglements in everyday settings align with previous work (Jansen et al., 2015; Frauenberger, 2019; Yu and Juan, 2025). Embedding sensing into the cushion promised low additional burden and minimal change to their physical environment.
4.1.2. Creating a Meaning-Making through Atmosphere (Phase 2).
Because connecting the choice of material and embodiment to interactions can help tailor a physicalization to a particular target audience (Li et al., 2025; Miao et al., 2025), we refined the concept into a two-part hybrid system: a sensing cushion plus an external display.
Initially, we tried numeric and bar chart visualizations, assuming they would support clear self-monitoring. But in the poster feedback from both older adults and experts revealed strong negative reactions: graphs felt clinical, judgmental, and reminiscent of hospitals, raising concerns about data anxiety and long‑term engagement. This pushed us towards ambient, non‑numeric feedback using a small, furniture‑like object inspired by (Bae et al., 2022; Suzuki et al., 2017).
@@ 4.1.3. Leveraging Abstraction and Feelings of Companionship for Meaning Making (Phase 3)
This sense of companionship was achieved by changing the room’s feeling. Participants perceived the device not as an information display but as felt it could integrate part of their home’s ambiance, providing a canvas upon which they could construct personal significance. P9_\_Phase3 said “It is not data, it is like an atmosphere, like when the sun sets, and the room gets warm and golden, you know the day is ending. This light is like that. It doesn’t count on me. It just changes the feeling of the room. This change in lighting is very romantic … a very gentle way. Its a very poetic way of knowing time has passed.”. By comparing the light to a sunset, the participant was not merely having an experience, he was actively interpreting the abstract data by connecting it to a familiar, natural phenomenon. This act of drawing an analogy transformed the device from an information display into a part of the room’s narrative. This demonstrates that for this user group, data accessibility is contingent not only on simplification but also on its integration into a form that allows for personal meaning-making.
@@ 5. Discussion
Our study aimed to explore how to make sedentary data accessible for older adults and how this method differs from traditional reminders. The results show that the key to accessibility is not to simplify quantitative data but to transform it into an atmospheric abstraction (Smith et al., 2025; Heijboer et al., 2016). Participants reported experiencing the light as gentle, non-clinical, friendly, and as an ordinary lamp that did not make them feel judged or subjected to a medical check-up. This is in line with prior work that argues for supportive instead of paternalistic health technologies for older adults (Yu and Juan, 2025; Voinea et al., 2024). In this sense, our findings showing that supportive and atmospheric abstractions can be maintained in everyday in-situ use, rather than only in short-term or laboratory settings. Herein, we reflect on our findings and describe their implications for future interventions that support older adults living, and interacting with, their personal data.
@@ 5.1. Rethinking Forms of Physical Data Representations for Older Adults
In light of our results, it is necessary to shift the model of interaction from one of command and compliance to one of suggestion and negotiation, which better protects a user’s sense of autonomy. Because the system did not provide precise numbers, it avoided generating feelings of being judged and monitored. Participants thus tended to perceive the device as a gentle and companion-like presence (Liu et al., 2023). Such positive emotional connections are difficult for traditional health tracking devices to establish, even though they can make health interventions feel more humane and could increase the likelihood they are accepted by users in the long term.
The ambiguity and gentleness of such designs may have potential negative impacts as well. If a design is too focused on being non-intrusive, it may struggle to achieve its original health intervention goals (Frissen et al., 2024; Ntsweng et al., 2025; Lee et al., 2023). When users ignore a reminder, it is also difficult to determine if this was due to an empowered, autonomous choice or if a design had failed. Rather than trying to find a perfect balance between clarity and ambiguity, it might be more prudent to actively manage this uncertainty and explicitly give final control to the user (Berridge et al., 2022). Our mute button allowed users to negotiate with the system and acknowledge a reminder while choosing to act on it later. Allowing users to negotiate, delay, or selectively decline prompts may help reduce feelings of being controlled or judged and may lower reactance and data-related anxiety (Thudt et al., 2018; Ramesh, 2024). It can also help preserve a sense of autonomy and dignity, which prior work has identified as a central concern in the design of technologies for aging populations (Li et al., 2025; Lupton, 2017). In line with the exploratory, in-situ nature of our RtD process, we suggest that such negotiated engagement should be seen as an informative outcome that complements more conventional measures of short-term effectiveness.
@@ 5.2. Leveraging Research Through Design for Exploratory Digital-to-Physical Translations
This experience thus suggests that for this class of problem, RtD can act as a form of experiential inquiry. As our central question was phenomenal rather than technical (i.e., what does it feel like to live with an ambient representation of your own data?), we needed to build artifacts and have end users live with them (Gaver et al., 1999; Lazar et al., 2010). Through this process, we uncovered the direct conflict between a desire for calm ambiance and the need for a clear, actionable alert (Lee et al., 2023; Pousman and Stasko, 2006). Our artifacts became physical arguments, each one testing a different stance on this conflict.
5.3. Living with Data through Ambient Co-dwelling
Our investigation suggests that making sedentary data accessible for older adults requires more than merely simplifying numerical visualizations. It demands a fundamental shift in the user-information relationship—a concept we frame as living with data. To live with data implies that personal health information is no longer treated as an external metric requiring active, focused interrogation. Instead, it becomes an integral, unobtrusive quality of the domestic environment (Kaziunas et al., 2018). In this paradigm, the goal shifts from the rigorous tracking of performance to the cultivation of a sustained awareness(VandeWeerd et al., 2020; van Hoof et al., 2011), where data is woven into the fabric of daily life rather than standing apart as a distinct object of scrutiny.
This state of living with data is realized through the mechanism of ambient co-dwelling, where quantitative data is transformed into an atmospheric abstraction (Smith et al., 2025; Ntsweng et al., 2025). Unlike traditional trackers that participants described as ticking death-countdowns (P12), the ambient light allowed data to dwell within the home as a gentle and friendly presence, akin to a sunset or a household pet. By integrating information into the feeling of the room (P9), the system shifts the experience from surveillance to co-dwelling. This approach enables users to engage with their health status through peripheral perception (Weiser and Brown, 1996; Heijboer et al., 2016; Bakker et al., 2016), allowing the data to exist comfortably in the background without demanding immediate attention or inducing judgment.
However, this prioritization of ambient comfort reveals a complex tension between living with data and the mechanics of joint action motivation (Hommes and Kipp, 2024). While the physicalization of sedentary behavior could mitigate the anxiety associated with surveillance, the resulting abstraction could inadvertently obscure the precise behavioral cues required to trigger synchronous social interventions (Dourish, 2001; Schmidt, 2002). For older adults in everyday life, the very features that make the system unobtrusive simultaneously dampen the immediate social pressure to act together (Rogers, 2006; Kotary and Hoyer, 1995). Consequently, physicalization approaches might navigate a delicate trade-off where the aesthetic integration of health information fosters long-term acceptance but may reduce the explicit urgency typically needed to drive mutual adherence and immediate behavioral change.
@@ 5.4. Design Implications
These implications outline a design direction for future tangible health technologies that shifts the emphasis from monitoring and compliance to empowerment, allowing older adults to maintain dignity and autonomy while data are woven into their everyday lives. In the context of health support, enabling older adults to reduce reactance, avoid data anxiety, and make it more likely that the intervention will be lived with over time.
@@ 5.5. Limitations and Future Work
While providing novel insights, our research does have several limitations. First, the one-day deployment of the CushLumi system during Phase 3 prevented us from understanding the potential long-term effects of the intervention, such as what occurs when its novelty wears off or what would occur if it were utilized outside a controlled, in-home setting. Our future plans include a longitudinal, in-situ study lasting several months to examine whether and how the atmospheric abstraction supports sustained changes in sedentary patterns and ongoing engagement over time. For instance, we are interested in understanding the degree to which atmospheric abstraction can lead to sustained behaviour change, and how users’ interpretations of the ambient light shift with prolonged use. It is important to note that evaluating the clinical efficacy or long-term health outcomes of the device was outside the scope of this exploratory design study, which focused on user experience and acceptance.
The small sample size of five older adults in Phase 3, also limits the generalizability of our findings to the broader older adult population and to other cultural or geographic contexts. As our research was conducted with older adults in an Australian Western context, future work must examine how ambient data physicalizations are perceived in other cultural settings, including non-Western contexts that may have different norms around aging, health, family dynamics, and family care.
Objective studies of sedentary behavior also typically rely on accelerometers within wearable devices to measure sedentary time (Healy et al., 2008b, a). Our cushion-based sensing plays a similar role, but within a piece of decor. We did consider the use of load cells, pressure mats, and camera-based sensing to measure sedentary time, however, load cells and mats required more rigid structures and careful calibration, while cameras raised privacy concerns. Flexible sensors offered a good trade-off between simplicity, comfort, and ease of replication. It would, however, be worthwhile to explore other sensing methods and form factors in the future.
Architecturally, the CushLumi system was also designed so that participants’ own sensor data could serve as a basis for secondary modelling and personalisation. This adaptive layer was not activated in the present study but will be explored in future work to better understand the well-being data. Furthermore, we plan to iterate on the CushLumi system’s design by exploring different materials, forms, and sensory modalities (e.g., subtle temperature changes) to expand the vocabulary of atmospheric physicalizations that are possible.
@@ 6. Conclusion
Using a three-phase Research through Design methodology, our research explored how sedentary data could be translated into an atmospheric, ambient physicalization for older adults. By offering gentle, suggestive guidance rather than intrusive commands, participants’ data anxiety was mitigated and older adults felt empowered to be less sedentary. Our CushLumi System prototype fostered positive, relational engagement with one’s personal health data, marking a shift from merely informing users to enabling them to live reflectively with their data.
Acknowledgements.
This research was undertaken while I was a student at The University of Queensland, and I gratefully acknowledge the support and facilities provided by the university. I also extend my sincere thanks to the older adults who participated in this study for their valuable contributions.
@@ Appendix A Design Explorations and Rationale via RtD
We initially explored the use of a smartphone application that aimed to deliver sedentary behavior feedback using conventional interaction methods and visualizations. During Phase 1, however, participants had an aversion to complex interfaces, small text, and the cognitive load of managing yet another app. For example, one participant remarked, “If I can choose, I don’t want to use smart phone, it has too many functions” (P8). Participants also explicitly stated that they wanted to avoid reminders that felt like nagging. This feedback, in addition to prior work that identified that longer sedentary periods are associated with higher clustered metabolic risk and all‑cause and cardiovascular mortality (Owen et al., 2010; Healy et al., 2008b, a) and recommendations that emphasized interrupting sitting every 30 to 60 minutes (Owen et al., 2010; Healy et al., 2008a; Martin et al., 2015; Prince et al., 2014), led to the exploration of interventions that could be embedded within existing physical environments and daily routines to discourage sedentary behavior.
@@
Building on the feedback obtained during Phase 2, we then designed and implemented the CushLumi System (Figure 2). Our design included a cushion and a light because prior ambient display work leveraged light for slowly changing, low priority information(Wisneski et al., 1998; Davis et al., 2017a; McLaughlin and Pak, 2020; Feine, 2025) and research on aging in place environments highlighted furniture (e.g., lighting) as a key interior design factor for supporting sleep, mood, and daily functioning in older adults (Yu and Juan, 2025).
To make the connection between the formative findings and the final form factors explicit, we summarize our design explorations across the three RtD phases in Table 2. For each phase, we articulate the design goals, the rationale grounded in user and expert feedback, the configuration we chose to pursue next, and representative participant quotes that directly motivated these choices.
Appendix B CushLumi System Implementation
The Sensing Cushion contained two Spectra Symbol 4.5-inch flexible sensors that were embedded in the foam of a 30 cm × 60 cm foam cushion. The sensors formed a shallow V pattern that was intended to roughly align with a user’s thighs and pelvis. Each flex sensor was sampled by a Particle Photon microcontroller at 10 Hz. The microcontroller read both analog channels, detected bending due to body weight, and aggregated readings over a 3 second window with hysteresis.
The Ambient Light was a small wooden box that featured laser-cut patterns of positive activities (e.g., stretching, gardening) on its faces, giving it the appearance of a tabletop lamp. The lamp measured approximately 8 cm × 8 cm × 8 cm and was powered by a single-cell LiPo battery, allowing it to be freely placed on side tables or shelves. The box contained 16 addressable WS2812 RGB LED Circle that glowed depending on one’s sitting time (i.e., 0 - 20 minutes -¿ no intervention needed so the LEDs were warm white, at 20 - 30 percent brightness; 20 - 40 minutes -¿ gentle prompting where the hue linearly interpolated from warm white to orange/red as a function of the elapsed minutes, with the brightness increasing moderately over time; 40+ minutes -¿ standing is recommended so the hue shifted toward saturated red, with the brightness increased to 70 to 80 percent, with smaller additional increases if the user did not stand up). In line with work showing that even replacing small amounts of sedentary time with light intensity movement can benefit metabolic health (Healy et al., 2008b, a), the mapping was intended to gently nudge users toward more frequent, small breaks instead of signaling a sudden point where it was already too late to see benefits of standing. The intensity levels conservatively operationalized the 30 to 60 minute recommendations reported in sedentary behavior research (Owen et al., 2010; Healy et al., 2008a; Dempsey et al., 2020). When a user stood up, a short celebratory animation was played (i.e., the lamp cycled through rainbow hues for 2 to 3 seconds). The lamp’s LEDs were updated at a frequency of 1 Hz. A touch-sensitive area on the top served as a Mute button, providing direct control over the feedback and reinforcing the user’s sense of agency. Five CushLumi Systems were built for use in Phase 3.
@@ Appendix B CushLumi System Implementation
cushion as a form factor
Understand older adults’ sedentary routines and attitudes to technology; explore initial ways to surface sitting time without disrupting everyday practices. Through observations and interviews, we found that many participants struggled with smartphones and avoided “one more digital device”. We also found that cushions repeatedly appeared as mundane but central objects in long sitting episodes (living rooms, bedrooms), which have opportunities to highlight the intelligence of a particular data-physicalized thing that can sustain intimate entanglements in everyday settings align with previous work (Jansen et al., 2015; Frauenberger, 2019; Yu and Juan, 2025). Embedding sensing into the cushion promised low additional burden and minimal change to their physical environment. “If I can choose, I don’t want to use smart phone, it has too many functions.” (P8) [On cushions] “This is something I always use when I sit here.” (P1, field notes)
@@
Clarify how sedentary data should be represented; test reactions to explicit quantitative displays; maintain comfort by separating sensing from display. Because connecting the choice of material and embodiment to interactions can help tailor a physicalization to a particular target audience (Li et al., 2025; Miao et al., 2025), we refined the concept into a two-part hybrid system: a sensing cushion plus an external display. Initially, we tried numeric and bar-chart visualizations, assuming they would support clear self-monitoring. But in the poster feedback from both older adults and experts revealed strong negative reactions: graphs felt clinical, judgmental, and reminiscent of hospitals, raising concerns about data anxiety and long-term engagement. This pushed us towards ambient, non-numeric feedback using a small, furniture-like object inspired by (Bae et al., 2022; Suzuki et al., 2017). “It reminded me of how I felt when I needed to go to the hospital.” (P15, on bar charts) “Corny”, “boring”, and “annoying”. (P15, on group discussion) “I feel a bit oppressed by it.” (P13, on explicit numeric feedback) It’s not just telling me good job for standing, it’s reminding me of something I love.(P11, on explicit numeric feedback)
@@
Realise a fully functional system that embodies an “ambient, gentle” data narrative; keep sensing invisible while offering legible but non-intrusive feedback; support user agency and positive associations with movement. In order to build the composite relation with a data-physicalized thing in terms of non-neutrality, mutual shaping, and shared intentionalities, we implemented the CushLumi system: the cushion stays visually unchanged and embeds flex sensors plus a microcontroller to detect sitting (Zhong et al., 2025). We aim to build and deploy the technological data-physicalized thing as it has intentionality in assembling the cohabiting human beings and other things in the situated context of everyday (Wakkary, 2021). At the same time, we also hope that through the cube’s form design, we can offer target users more activity suggestions after standing, and motivate users to stand up proactively through intuitive visual cues (Anderson, 2011). A small wooden “ambient light” box sits on nearby furniture and gradually brightens as sitting accumulates; when the person stands, it shifts to a pleasant colour shining through laser-cut patterns of everyday activities (e.g., stretching, gardening), gently suggesting alternatives to prolonged sitting. A touch-sensitive area lets users mute feedback, reinforcing a sense of control. Participants often described the box as “a little lamp” or “a decoration”, not as a life countdown reminder. “I didn’t feel pushed. I just thought, ‘Maybe I’ll water the plants.’ And I did.” (P4, deployment interview)